Sandy Brondello scored 24 points in a recent basketball game without making any 3-point
shots. She scored 16 times, making several free throws worth 1 point each and several field
goals worth two points each. How many free throws did she make? How many 2-point
field goals did she make?

Answer by ikleyn(52787)

.
Sandy Brondello scored 24 points in a recent basketball game without making any 3-point
shots. She scored 16 times, making several free throws worth 1 point each and several field
goals worth two points each. How many free throws did she make? How many 2-point
field goals did she make?
~~~~~~~~~~~~~~~~~

Let T be the number of 2-point field goals;

then the number of the 1-point throws is (16-T).


Write the total score equation

    (16-T) + 2T = 24.


From the equation, find T

    16 - T + 2T = 24

           T    = 24 - 16 = 8.


ANSWER.  8 2-point shots and 24-16 = 8 free throws.

Solved.